OpenStudy (anonymous):

Let θ be an angle in quadrant 3 such that secθ= -(5)/(4) Find the exact values of cotθ and sinθ

OpenStudy (cwrw238):



OpenStudy (anonymous):

in quadrant 3, tan is +ve while sin and cos are -ve sec x=(-5/4) cosx = -4/5  y=sqrt(25-16)=3 sinx=-3/5 cotx=cos/sin =.... secx = -5/4 as it is given cscx =1/sin=.....